OpenStudy (anonymous):

4.Given that y=2x^2+3x, find the range of values of x for which

OpenStudy (anonymous):

\[y>9\]

OpenStudy (anonymous):

@hartnn @ganeshie8 @mathmath333

hartnn (hartnn):

did you first try yourself ?

hartnn (hartnn):

its the same as last one, give it a try, show your steps

OpenStudy (anonymous):

i did try @hartnn

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[9=2x^2+3x\] \[-2x^2-3x+9=0\] \[2x^2+3x-9=0\]

hartnn (hartnn):

good how will you factorize that ?

hartnn (hartnn):

nopes, find 2 numbers whose sum is +3 and product is -18

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

\[-3\times6=\]

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

\[-3+6=3\]

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

correct so now we can factorize \(2x^2+3x-9 = 2x^2 +6x-3x-9 \\ = 2x (x+3)-3(x+3)\) can u continue ?

hartnn (hartnn):

factor out x+3

OpenStudy (anonymous):

x=-3 and x=3/2

hartnn (hartnn):

yeah, but keep it in factored form

hartnn (hartnn):

(x+3)(2x-3)

OpenStudy (anonymous):

okay

hartnn (hartnn):

\((x+3)(2x-3) > 0 \) how will u go ahead ?

OpenStudy (anonymous):

\[x >-3~and~x >\frac{ 3 }{ 2 }\]

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

thnx @hartnn